Circular No. 8272

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004G IN NGC 5668 S. Nakano, Sumoto, Japan, reports the discovery by R. Kushida (Yat-

sugatake South Base Observatory) of an apparent supernova (mag 17.2) on numerous unfiltered CCD frames taken around Jan. 19.86 UT with a 0.40-m f/10 Schmidt-Cassegrain reflector. Y. Kushida provides the following measurement for the new object: α = 14^{h}33^{m}21^{s}.40, δ = +4^{o}26′49′′.5 (equinox 2000.0), which is 43′′ west and 12′′.5 south of the center of NGC 5668. Nothing was visible on numerous past frames taken by R. Kushida, the most recent having been made on 2003 Apr. 8 (limiting mag 18). A confirming image was obtained by K. Itagaki (Teppo-cho, Yamagata, Japan) on Jan. 20.722, showing SN 2004G at unfiltered mag 17.5 and at position end figures 21^{s}.39, 50′′.1. Itagaki also found a prediscovery image of the new object at mag 16.8 on Jan. 11.78; nothing was visible on an image taken by Itagaki on 2001 May 28 (limiting mag 18.5). Supernovae 1952G and 1954B also occurred in this galaxy.

SUPERNOVAE 2004H, 2004I, AND 2004J Further to IAUC 8271, B. Swift, J. Burket, J. Graham, R. J. Foley, and

W. Li report the LOSS discovery of three apparent supernovae on unfiltered KAIT images:

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004H Jan. 17.5 11^{h}33^{m}59.^{s}77 +49^{o}03′46′′.5 17.6 5′′.9 E, 3′′.5 N
2004I Jan. 18.1 2 43 31.23 + 0 18 30.7 17.2 1′′.1 W, 6′′.4 N
2004J Jan. 19.2 5 46 32.18 −17 46 53.0 17.7 8′′.5 W, 1′′.8 N

Additional unfiltered KAIT magnitudes: SN 2004H in IC 708, 2004 Jan. 2.5 UT, [18.5; Jan. 19.5, 17.6. SN 2004I in NGC 1072, 2003 Nov. 21.1, [19.0; 2004 Jan. 19.1, 17.1. SN 2004J in ESO 554-G33, 2003 Nov. 18.2, [18.5; 2004 Jan. 20.3, 17.5.

POSSIBLE NOVA IN NGC 205 Further to IAUC 8267, J. D. Neill reports the discovery on CCD images

of a possible nova (mag V = 19.6 on Jan. 17.146 UT and V = 19.7 on Jan. 18.141) located at α = 0^{h}40^{m}15^{s}.24, δ = +41^{o}37′29′′.9 (equinox 2000.0), which is ≈ 76′′.9 west and 217′′.2 south of the center of NGC 205. The new object is present on images back to 2003 Dec. 25, but not on images taken on Dec. 21 (limiting mag 20.5) and earlier or on Palomar Sky Survey plates.

2004 January 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green